OpenStudy (anonymous):

Find the sum of the summation of 3 i plus2, from i equals 5 to 13.

OpenStudy (tkhunny):

Have you considered substituting the values given for 'i' and adding them up? Are you supposed to use the Arithmetic Series sum formula? Show your work!

OpenStudy (anonymous):

Sn=n/2(a1+an)?

OpenStudy (tkhunny):

I'll take that as a 'yes' to the second question. Okay, how do you interpret the formula, in light of the problem statement? What is a1? What is n? What is an?

OpenStudy (anonymous):

a1=17

OpenStudy (anonymous):

n=13

OpenStudy (anonymous):

an is the number youre looking for

OpenStudy (tkhunny):

a1 = 3*5 + 2 = 15 + 2 = 17 -- Good. n = count(5, 6, 7, 8, 9, 10, 11, 12, 13) = 9 -- Whoops!

OpenStudy (anonymous):

okay! so thats where i was going wrong

OpenStudy (tkhunny):

n is the number of terms. Never underestimate the power of simply counting.

OpenStudy (anonymous):

is the answer 261?

OpenStudy (tkhunny):

You tell me.

OpenStudy (anonymous):